OpenStudy (anonymous):

Use the point-slope form of the linear equation to find an equation of the line with the given slope and passing through the given point. Then write the equation in standard form. Slope 1/4 ; through (-3, 8) a. x - 4y = -44 b. x - 4y = 35 c. x - 4y = -35 d. 4x - 4y = -35

OpenStudy (angela210793):

same logic as before Lijustine

OpenStudy (anonymous):

well, you do (y-y1)=m(x-x1) which is y-8=1/4 (x+3) mult both sides by 4, so 4y-32= x+3. subtract 3 from both sides and subtract y from both sides. so the answer is C